### v1

# Fluidic FlowBots: Intelligence embodied in the characteristics of recirculating fluid flow

Abstract
The one-to-one mapping of control inputs to actuator outputs results in elaborate routing architectures that limit how complex fluidic soft robot behaviours can currently become. Embodied intelligence can be used as a tool to counteract this phenomenon. Control functionality can be embedded directly into actuators by leveraging the characteristics of fluid flow phenomena. Whilst prior soft robotics work has focused exclusively on actuators operating in a state of transient/no flow (constant pressure), or pulsatile/alternating flow, our work begins to explore the possibilities granted by operating in the closed-loop flow recirculation regime. Here we introduce the concept of FlowBots: soft robots that utilise the characteristics of continuous fluid flow to enable the embodiment of complex control functionality directly into the structure of the robot. FlowBots have robust, integrated, no-moving-part control systems, and these architectures enable: monolithic additive manufacturing methods, rapid prototyping, greater sustainability, and an expansive range of applications. Based on three FlowBot examples: a bidirectional actuator, a gripper, and a quadruped swimmer - we demonstrate how the characteristics of flow recirculation contribute to simplifications in fluidic analogue control architectures. We conclude by outlining our design and rapid prototyping methodology to empower others in the field to explore this new, emerging design field, and design their own FlowBots.
I Introduction
I-A Embodied intelligence expands the capabilities of fluidic soft robots.
The field of robotics has seen a plethora of developments in using fluid pressure (pneumatic and hydraulic) to actuate compliant, soft-bodied robots [1]; these soft-robots exhibit superior capabilities over their rigid counterparts in use cases requiring high dexterity and safety during operation with humans [2, 3]. Replacing electronic controls with fluidics has also opened up the possibility of deploying soft robots in extreme environments, thanks to the inherent high power density, and temperature and radiation resistance of fluidic systems [4, 5, 6].
There have been a range of notable advancements in recent years for controlling fluidic soft robots using hardware that mimics digital electronic devices and architectures [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Despite these advancements, scaling fluidic control system to higher degrees of complexity remains difficult; a large number of pressure control inputs is required, and routing control signals becomes an issue [4, 23].
Attempts at overcoming this “routing problem” have included fluidic demultiplexer chips [15], oscillators [13], fluidic matrix circuits [23], as well as analogue devices that forego the need for discretisation [14, 24]. An alternative approach involves embodied intelligence, whereby control features can be embedded directly into the robot. By exploiting the mechanics of the interaction of the robot with its physical operating environment as a form of sensory feedback, the need for external control inputs can be reduced [25, 26, 27, 28].
I-B Internal flow characteristics can embed functionality into the structure of the actuator.
In most reported-cases, and for each of the actuators cited by Section I-A, the pressure is adjusted by pumping set quantities of fluid between the individual chambers of the actuator; this transition is carried out in a one-off, pulsatile, or cyclic manner. Due to the mechanical coupling between the fluid medium and the vessels/channels containing it, the behaviour of a fluidic actuator can be directly influenced by controlling the pressure of the fluid inside of it. During the pumping phase, the flow rate of the fluid is determined by its interaction with the flow channels; this mechanical coupling influences the rate of change of the pressure distribution inside the actuator, and, therefore, the deformation of the actuator. If the characteristics of the flow can be exploited, they can be used to control the actuator.
Viscous energy losses lead to a decrease in the pressure of the flowing fluid. Previous work in the field has explored exploiting this phenomenon for control purposes [29, 30, 31, 32, 33]. In their recent work [34], Matia et al. used the resultant pressure asymmetries to control the gait of a walking robot. Matia et al. accurately modelled the relationship between the timing of viscous phenomena and the inertial coupling between the fluid and the actuator; therefore, they demonstrated how controlling the timing of transient fluid flows can serve as a form of embodied intelligence.
Whilst there as several other notable examples of previous fluidic soft robots that used “pulsatile”, or “cyclic” modes of flow (such as [35, 8, 36, 37, 19, 38]), a condition that has not yet been explored and well-characterised in the context of soft fluidic robots is a state of continuous flow recirculation. Figure 1 illustrates the difference between soft fluidic robots based on the “steady state”, “pulsatile/cyclic”, and “recirculating” modes of flow, as defined in this paper. Fluid can be recirculated in a system for example to achieve thermal regulation, as in the work of Cacucciolo et al. [39], and others [40, 41]. Liao et al. [41] showed that the pressure of this recirculating fluid can also be used to inflate the cavities of an actuator, achieving deformation whilst simultaneously taking advantage of the thermal regulation characteristic (which they used to modify the stiffness of the actuator).
In this paper, we delve deeper into the behaviour of soft robots operating under the closed-loop fluid recirculatory condition, describing the opportunities granted by it in the way of embodied intelligence, as well as some of the associated design challenges. We aim to empower designers to build a whole new class of fluidic robots, which we term “FlowBots” (as further described by Section I-D).
Figure 1: Soft fluidic robots utilising a) ”steady state” b) ”pulsatile/cyclic” c) ”recirculating” flow, as defined in this paper.
I-C Intelligence embodied in fluid flow characteristics: from shark intestines and Tesla-valves to soft robots that can be printed as one part.
The state of “recirculation” is characterised by the driving fluid, taken from a reservoir, being pumped, through the channels in the actuator channels, and back to the reservoir, as illustrated on Figure 1. To use an electrical analogy, one can think of the actuator as a source of impedance in a closed-loop fluidic network. The ”current” (the driving fluid) flows from a ”high potential” (outlet of the pump/pressure source), through the ”impedance” (internal channels in the actuator), and back to a low potential (the fluid reservoir from which the fluid was originally pumped). Energy is lost via viscous friction, which constitutes the resistive component of the fluidic impedance, and the drop in fluid pressure that results is analogous to a drop in electrical potential.
There are also other fluid flow phenomena that can contribute to a change in the pressure distribution in the system. Some of these phenomena deal with the inertial component of the fluidic impedance, which is analogous to electrical inductance. A simple example is the Tesla-valve [42], where the geometry of the device results in pronounced impedance in one flow direction compared to the other. As visualised by Figure 2(a,b), the added secondary flow ”loops” impede the main flow when they collide with it, as a result of the inertia they carry. In nature, a similar geometry can be found in the intestinal tracts of certain species of sharks, skates, and rays. Leigh et al. [43] demonstrate that this geometry slows the flow in one direction due to the added fluidic impedance; they hypothesise that this geometry helps the animals slow the rate of food transit (thereby improving nutrient absorption in the short intestinal tract), as well as reducing the energy cost of digestion due to promoting flow in only one direction. The advantage of this type of geometry is that it avoids moving components (such as the flaps found in heart valves), which would be prone to blockage.
Figure 2: Inspiration for exploiting fluid flow phenomena as a form of embodied intelligence can be found in nature [43]. The geometry of the shark intestine (c) resembles that of a Tesla-valve (a, b). Flow circulation induced by the bellowed geometry results in a diode-like behaviour, making it harder for fluid to flow in one direction (left-to-right as pictured on the figure) than the other (right-to-left). This type of fluidic diode has no moving components; the added functionality is embodied into the structure of the device.
I-D FlowBots: a new class of soft robots with intelligence embedded in the characteristics of recirculating fluid flow.
This paper focuses on embedding additional functionality directly into the bodies of soft robots, by taking advantage of the properties of recirculating fluid flow. Here, we demonstrate that new types of soft robots can be manufactured as a single part, with embedded control, and actuation. This mode of manufacture enables rapid prototyping, increases robustness, reduces waste, and minimises the risk of environmental pollution (opening up opportunities for new applications in e.g. wildlife biodiversity monitoring). We term such robots “FlowBots”.
FlowBots are designed to operate under the state of recirculating flow, as defined by Section I-B; this characteristic differentiates them from previous soft robots controlled and actuated using fluids. Here, fluid flows into an inlet port and flows out through an outlet port. A single actuator can have multiple inlets and outlets, as well as additional control, or sensing ports if needed. FlowBots are capable of complex behaviours with only a single pressure source, due to their embodied fluidic intelligence.
The focus of this paper is to demonstrate the principles of operation of FlowBots based on three simple examples: a bidirectional actuator, a gripper, and a quadruped swimmer robot. Each example builds upon the operating principles of the previous one. The bidirectional actuator demonstrates how pressure asymmetries arising due to viscous losses can be used for analogue control using localised flow channel constrictions. The gripper shows how flow recirculation can simplify control architectures of multi-actuator systems due to the ability of switching between connection in parallel, and in series. The swimmer FlowBot builds upon the control architecture used for the gripper, and serves as a demonstration of how the embodiment of intelligence in fluid flow characteristics allows to create robots capable of complex behaviours, yet manufactured as single, robust components, without any moving parts. Section II describes some of our experimental work in evaluating the performance of FlowBots with air and water as the operating fluid. We conclude with a discussion on some of the design challenges associated with building FlowBots, empowering others in the field to build their own.
II Results
II-A Localised pressure asymmetries arising due to viscous losses in fluid flow can be used for analogue, bidirectional control.
Figure 3 illustrates the operation of the first FlowBot described in this paper – a bidirectional actuator. The left and right pressure chambers of the actuator are connected at the tip to allow for recirculation of flow from and back to the pressure source in a closed loop. Energy is lost due to the friction between the flowing fluid and the walls of the channel, and the pressure of the fluid progressively decreases as it advances along the flow path constituted by the internal channels inside the actuator. By concentrating these viscous energy losses at the tip of the actuator, a pressure distribution is obtained that is symmetric along the longitudinal centreline of the actuator; the chamber in one half has high pressure, and the other half has low pressure. This pressure asymmetry results in deformation in one direction. When the flow direction is reversed, the asymmetry flips, and the actuator deforms in the opposite direction.
If the location of the constriction was moved further upstream, or downstream of the tip, the deformation characteristic of the actuator would be permanently skewed to one side. Having the actuator deform differently in both directions at the same pressure difference could be desirable in certain applications. In this paper, all the actuators are symmetric, but we utilise external variable constrictions to regulate the direction of flow, and the pressure difference across the actuator. These constrictions can be implemented using additional valves, by manually blocking/kinking the outlet vents/tubes, or by using soft buttons (see the supplementary video). Figure 3 shows how we were able to use two such constrictions to achieve analogue control of the actuator in two directions, by regulating the pressure across it over a continuous spectrum. This analogue mapping can also be achieved by adjusting the direction of flow, as well as the flow rate of the supply pump (hydraulic case), or compressed air supply pressure (pneumatic case).
Figure 3: The principle of operation of the bidirectional actuator. By varying the outlet vent constrictions, the direction, and magnitude of the flow, as well as the pressure difference across the actuator can be controlled with an analogue characteristic. Reversing the flow direction causes the actuator to deform symmetrically in the other direction due to the mirroring of the pressure asymmetry. The percentage value of the constriction is an approximation; we estimated it indirectly by measuring the pressure difference across the actuator using two pressure gauge probes.
II-B FlowBots have simpler control architectures due to being able to operate in parallel, as well as in series.
Section II-A demonstrated how to control a single bidirectional actuator, two controls (vent outlet constrictions) are used. To fully control two such actuators, four controls would normally be needed. By exploiting the characteristics of recirculating fluid flow, we were able to reduce this fan-in to three. The simplification comes from the fact that under the state of recirculating flow, the actuators can operate not only in a parallel, but also a series connection. The ability to switch between the two configurations allows to simplify the control architecture, and opens up the design space for constructing FlowBots consisting of several actuators, that are capable of more complex behaviours.
Figure 4 illustrates the operation of the gripper that we have built using two bidirectional actuators described in Section II-A. The gripper uses only three control inputs rather than four, yet each of its fingers can still be controlled fully independently. The last two configurations are made possible by the fact that if the middle port is blocked, the two actuators become connected in series rather than in parallel. This capability would not be possible without using recirculating fluid flow in the system.
Figure 4: The principle of operation of a gripper consisting of two bidirectional actuators. By switching between a parallel ((a) to (g)) and series ((h) and (i)) connection of the actuators, we were able to achieve fully independent analogue control of each of the fingers using only three control inputs rather than four. Note that in the series configurations the actuators deform slightly less, due to the pressure difference across each one being effectively half of that in the parallel configuration.
II-C Complex FlowBots can be additively manufactured as one part, with no moving components.
The control characteristics of FlowBots are embodied in the interaction of the fluid flow, and the structure of the robot structure. No added parts, or moving components are required, and, as a result, entire robots can be additively manufactured as a single part, in one print job. Figure 5 depicts the operation of a quadruped swimming robot that we have built based on the control architecture used for the gripper design described in Section II-B. As in the case of the gripper, we can fully control each pair of limbs for this FlowBot using only three control inputs rather than four, thanks to the ability to switch between parallel and series connection of the actuators. Despite the added complexity of the design, the entire 3D stack of the control architecture comes packaged into the body of the robot, and has been additively manufactured as a single component. After the print job is over, the robot is ready to use. Embodying intelligence in this way makes it easy to prototype FlowBots with ever more complex 3D designs.
Figure 5: Operation of a quadruped swimmer FlowBot that builds upon the control architecture used for the gripper described by Section II-B. Due to the embodiment of control into the body of the robot, as well as the characteristics of recirculating flow, FlowBots of increasing complexity can still be manufactured as single components.
II-D FlowBots can operate with water, or air, as the working fluid.
The mechanical properties of the working fluid affect the characteristics of the recirculating flow. The fluid viscosity influences the flow rate in the system. The same flow rate with a more viscous fluid, such as water, will result in larger viscous losses, and therefore larger pressure drops than with a less viscous fluid, such as air. Nevertheless, for the same magnitude of pressure difference across an actuator, we expect the deformation to be identical, as the deformation of the bellows of the actuator is driven by the static component of the pressure. To validate this assumption, we conducted an experiment where we measured the deformation (curvature) of the bidirectional actuator described in Section II-A with air, and water, as the working fluid. Figure 6 shows the results of the experiment; these results are discussed in Section III-A.
To account for the difference in the viscosity of the fluids, we used the pressure difference between the inlets of the actuator as the independent variable in the experiment. This approach allowed us to validate whether the actuator deforms the same amount with the same pressure difference, independent of whether air, or water, was used as the working fluid. We examined a range of pressures from 1.25 bar to 2.5 bar (in intervals of 0.25 bar), testing in both flow directions with 3 measurement repetitions per data point. We also repeated the measurements with another identical actuator specimen. Due to the symmetry resulting from placing the channel constriction at the tip of the actuator, we expect it to deform the same amount in both directions, as explained in Section II-A. For more details about the experimental setup, please refer to the supplementary material.
Figure 6: The bidirectional actuator deforms consistently with both incompressible (water), as well as compressible (air) media as the working fluid. We attribute the observed variations between the measurements to micro-errors in the manufacturing process on the low-cost 3D printer that we used. Note that the error bars represent the deviation in measurements between the 2 actuator specimens that were tested. The error between the 3 measurement samples gathered during each of the experimental runs was negligible. The measured fluid flow rates at each pressure are annotated on the figure, although they were not explicitly needed for this specific analysis.
III Discussion
III-A FlowBots are compatible with both pneumatic, as well as hydraulic control.
Figure 6 shows the analogue response of the actuator with pressure with both air and water as the working fluid. We observe that, the deformation of the actuator is consistent between the two operating media in both directions within the examined pressure range. There was a variance in the deformation between the two directions of up to 11%; the examined specimens deform slightly more in the counterclockwise direction. Based on our experience manufacturing and testing different FlowBots, we have observed that this magnitude of error is common, and is related to micro-errors in the manufacturing process. One of the main factors at play appears to be bed levelling, which can cause the printer to extrude a different amount of material in certain locations. This phenomenon can be more pronounced on budget 3D printers such as the one used here. Due to the fact that the actuator samples used in this experiment were printed in the same location and orientation on the bed in two sequential print runs, it is possible that the bed was misaligned in one direction, or became misaligned during the print, and contributed to the error.
The channels in the actuators featured in the experiment are as narrow as 0.6 mm, and variations on the order of hundreds of microns at different locations along the flow path can disturb the symmetry between the pressure distributions in both directions; such deviations can contribute to marginal differences in the deformation in two directions. In our case, these errors did not significantly affect the robot’s desired functionality, but it might be a factor to consider for those designing FlowBots with very specific precision requirements. In these instances, we recommend carrying out adequate testing to ensure that the desired level of precision has been achieved.
Please refer to the supplementary material for a discussion of our experiences with mitigating other micro-sources of variance that we have observed when carrying out our experiments, such as the hygroscopic properties of the elastic TPU filament that we used to print the robots, temperature variations in the working fluid reservoir, or small internal, or external leaks in the membranes of the actuators.
In spite of the uncertainty associated with the detected errors in measurement, our results demonstrate that FlowBots can operate consistently using both compressible, as well as incompressible media (air and water) as the driving fluid. This characteristic of cross-compatibility gives greater flexibility to designers and makes it possible to deploy FlowBots in a wider range of operating scenarios.
III-B Recirculating flow, and facile 3D Printing enables a wide design-space and low-barrier to entry for researching and using FlowBots.
This paper demonstrates how designers can come up with entirely new fluidic soft robot concepts by taking advantage of the characteristics of fluid flow. Constant recirculation of the working fluid comes with the cost of energy expenditure, but it allows to simplify control architectures (as demonstrated in Section II-B), and to rapidly prototype and additively manufacture entirely new FlowBot designs. For additional design considerations to take into account when adapting systems to use recirculating flow, please refer to the supplementary material, where we have discussed these points in further detail.
Recirculating flow
The exact behaviour of the fluid inside of the internal channels is difficult to characterise analytically due to the turbulence arising as a result of the combination of bellow geometries, multi-directional pipe bends, and varying channel cross-sections. Whilst we do invite others specialising in the domain to help better characterise these phenomena, we encourage those interested in coming up with their own FlowBot designs to take advantage of Finite Element Method (FEM) software. In our design methodology, we have been able to utilise Computational Fluid Dynamics (CFD) to obtain an approximation of the internal flow field that was more than adequate enough to facilitate effective design exploration and validation. By combining CFD with structural simulations using Finite Element Analysis (FEA) software we were able to simultaneously explore how modifying parameters in the design affects the fluid flow in the control system, as well as the mechanical deformation of the actuators.
3D printing
With our chosen method of Fused Deposition Modelling (FDM) 3D printing, we were able to manufacture all of the FlowBots featured in this paper as one part, without any supports or additional post-processing needed. Once taken off the print bed, each FlowBot is ready to use, as it integrates actuation, control, and even fluidic fittings in its design. In spite of the complex bellow geometry of its actuators, and the multi-layered internal 3D control architecture, the swimmer FlowBot described in Section II-C took under 7 hours to print on a cheap, \$200 Artillery Sidewinder X1 printer using Recreus Filaflex 70A filament (as well as 82A for the specimens in Section II-D). This design approach is more environmentally sustainable, adds robustness and saves mass and space due to the reduced need for complex routing/manifolds that come with a pressurising/valving system.
A design variable that we have not yet explored, but which could help open up the design space even further is the observation that the fluid being recirculated throughout the system carries energy. This energy could be extracted for useful work, for example by integrating rotary tool end effectors into the robots. We foresee this could be an interesting branch to explore for those interested in developing novel FlowBot systems, in particular ones oriented toward more specific industrial applications.
IV Conclusion
This paper introduced the concept of FlowBots, a new generation of soft robots that operate under a state of fluid flow recirculation. FlowBots are controlled by taking advantage of fluid flow characteristics, allowing for embodied intelligence. This work demonstrates the operation of three simple FlowBot systems: a bidirectional actuator, a gripper, and a quadruped swimmer robot.
By utilising pressure asymmetries arising from viscous losses during fluid flow, in combination with the ability to operate in series, as well as in parallel - features only granted by flow recirculation - this work shows that we have been able to simplify the control architectures required to achieve fully independent, analogue, and bidirectional control of all of the end effectors.
Each FlowBot was additively manufactured as a monolithic part with no moving components, making it low-cost, robust, and easier to dispose of at the end of its life cycle than multi-part systems utilising different types of materials. This paper outlines an efficient design methodology that will allow engineers to rapidly prototype their own FlowBots using FEM methods, and additive manufacturing.
This paper lays the foundations for a new and wide design space for soft robotic systems that use recirculating flow.